OpenStudy (thecuteone25):

Will MEDAL –2(n – 4) = 6 * Solve for n

OpenStudy (thecuteone25):

@jadzia @pooja195

pooja195 (pooja195):

i think you're missing parts of your question can you post the full question

OpenStudy (thecuteone25):

that's the whole question but I will post it again for you

OpenStudy (thecuteone25):

10.) –2(n – 4) = 6 * Solve for n

OpenStudy (anonymous):

HELP ME PPOJA

OpenStudy (thecuteone25):

the only thing new is the 10 but that means it's number 10

OpenStudy (anonymous):

POOJA

OpenStudy (anonymous):

No one will help me :(

pooja195 (pooja195):

ok start by distributing 

OpenStudy (thecuteone25):

ok so you have to subtract -2 on both sides?

pooja195 (pooja195):

Nope we multiply using the distributive property. a(c+b) = ac*ab So... -2 * n =? -4*-2=?

pooja195 (pooja195):

there's more to the process :P Can you answer : -2 * n =? -4*-2=?

OpenStudy (thecuteone25):

-6

OpenStudy (thecuteone25):

wait N=-4

OpenStudy (thecuteone25):

so both equations = -6

pooja195 (pooja195):

O_o I am so confused. –2(n – 4) = 6 * * <---is the supposed to be there?

OpenStudy (thecuteone25):

that sign is not there

OpenStudy (thecuteone25):

there are parentheses

pooja195 (pooja195):

–2(n – 4) = 6 so just this?

OpenStudy (thecuteone25):

yes

OpenStudy (mathstudent55):

Don't distribute the -2. Start by dividing both sides by -2.

pooja195 (pooja195):

oh ok –2(n – 4) = 6 when we distribute we get -2n+8 -2 * -4=8 (remeber that a negative times a negative is a positive ) -2+8=6 @mathstudent55 either way would work.

pooja195 (pooja195):

-2n+8=6 subtract 8 from both sides. 8-8=0 6-8=?

OpenStudy (thecuteone25):

-2

pooja195 (pooja195):

good -2n=-2 divide both sides by -2 -2/-2=1 so you have n -2/-2=? (on the other side)

OpenStudy (thecuteone25):

1

pooja195 (pooja195):

Good so n=?

OpenStudy (thecuteone25):

1

pooja195 (pooja195):

Perfect :)

OpenStudy (thecuteone25):

thanks

pooja195 (pooja195):

yw ^_^